# Learning thermodynamically constrained equations of state with uncertainty

### v1

@@ Abstract
Numerical simulations of high energy-density experiments require equation of state (EOS) models that relate a material’s thermodynamic state variables – specifically pressure, volume/density, energy, and temperature. EOS models are typically constructed using a semi-empirical parametric methodology, which assumes a physics-informed functional form with many tunable parameters calibrated using experimental/simulation data. Since there are inherent uncertainties in the calibration data (parametric uncertainty) and the assumed functional EOS form (model uncertainty), it is essential to perform uncertainty quantification (UQ) to improve confidence in the EOS predictions. Model uncertainty is challenging for UQ studies since it requires exploring the space of all possible physically consistent functional forms. Thus, it is often neglected in favor of parametric uncertainty, which is easier to quantify without violating thermodynamic laws. This work presents a data-driven machine learning approach to constructing EOS models that naturally captures model uncertainty while satisfying the necessary thermodynamic consistency and stability constraints. We propose a novel framework based on physics-informed Gaussian process regression (GPR) that automatically captures total uncertainty in the EOS and can be jointly trained on both simulation and experimental data sources. A GPR model for the shock Hugoniot is derived and its uncertainties are quantified using the proposed framework. We apply the proposed model to learn the EOS for the diamond solid state of carbon, using both density functional theory data and experimental shock Hugoniot data to train the model and show that the prediction uncertainty reduces by considering the thermodynamic constraints.
@@ I Introduction
EOS models are typically constructed using semi-empirical functions where the functional form is motivated by the physics, and the parameters are calibrated using a complex combination of experimental and first-principles simulation data from a variety of sources. Once calibrated, the EOS model can be used to interpolate and extrapolate over the wide range of input states needed in hydrodynamic simulations. The semi-empirical approach is subject to two sources of uncertainty; uncertainty in the values of the parameters in the EOS model (parameter uncertainty) and uncertainty in the form of the EOS model itself (model uncertainty). Both sources must be quantified simultaneously to give a complete picture of the total EOS uncertainty. While parametric uncertainty in the EOS has been addressed in several recent works, model uncertainty remains a significant challenge. In this work, we describe a new machine learning based approach to UQ which accounts for all sources of uncertainty and provides an analytical framework for combining heterogeneous data sources into a single, uncertainty-aware EOS model.
@@ II.1 Relevant Prior Work
In the standard setting, EOS parameter calibration depends on individual modelers who leverage domain knowledge and expertise to align EOS predictions with given experimental and simulation data. Recently, it has been common to pose the calibration process as an optimization problem to solve for the best parameters that give the least prediction error compared to available dataForte et al. (2018); Bergh, Wedberg, and Lundgren (2018); Cox and Christie (2015); Myint et al. (2021). This approach can naturally be extended to capture parametric uncertainty by considering uncertainties in the calibration data sets. Ali et al. (2020) proposed a method that considers small perturbations in experimental data to calibrate model parameters using an optimization routine and propagate the experimental uncertainties using Monte Carlo simulation through the EOS models. Brown and Hund (2018) apply Bayesian model calibration to estimate parameters using dynamic material properties experiments under extreme conditions. Lindquist and Jadrich (2022) proposed a Bayesian framework to perform UQ of a multi-phase EOS model for carbon by accounting for calibration data uncertainty and yielding an ensemble of model parameters set. Walters et al. (2018) uses a Bayesian statistical approach to quantify parametric uncertainty by coupling hydrocode simulations and velocimetry measurements of a series of plate impact experiments of an aluminum alloy.
@@ II.2 Uncertainty in Parametric EOS
UQ for existing parametric EOS models is further limited by the prescribed form of these models. Although often derived from physical principles, these models are nonetheless built upon assumptions, simplifications, and approximations of known physics, while neglecting physics that are poorly understood. Consequently, these models may be very accurate in certain regimes (e.g. of $T$ and $P$) and inadequate in others. The resulting uncertainty in these predictions is referred to as model-form uncertainty and cannot be accounted for in existing parametric models. In certain cases, competing parametric models can be compared and selected using Bayesian model selection. However, this requires the computation of the evidence term (denominator in Eq. (3)), which poses significant practical challenges. Model-form uncertainty, combined with parametric uncertainty, results in a range of outputs for a fixed input state, thus yielding an ensemble of valid EOS models. Hence, it is also necessary to quantify these model-form uncertainties in a rigorous UQ framework to enhance our confidence in EOS predictions. Significant research efforts have been made in quantifying parametric uncertainty; however, model-form UQ for equations of state has not received adequate attention, and only a few publications are available in the literature Kamga et al. (2014); Gaffney, Yang, and Ali (2022). The biggest hurdle in quantifying model-form uncertainty is enumerating all the potential mappings that could form physics-consistent EOSs. In this work, we develop a framework using Gaussian process regression that automatically explores the range of physics-consistent EOS models to capture both sources of uncertainty. The proposed method is non-parametric and data-driven, yet satisfies both thermodynamic consistency and stability constraints.
@@ II.3 Parametric EOS with Uncertainty: An Illustration
$\theta_{D}=\theta_{0}x^{-1.5}\exp[\gamma_{1}(1-x^{3q})/q]$ (7)
@@
Fitting the model therefore requires a complicated calibration process to infer the following vector of six parameters $\alpha=\{V_{0},K_{0},K^{\prime}_{0},\theta_{0},\gamma_{1},q\}$ from data at various pressures and temperatures Dewaele et al. (2008); Occelli, Loubeyre, and LeToullec (2003). This calibration has been performed in the literature Dewaele et al. (2008) and we have further conducted a preliminary Bayesian parameter estimation using the same data with the resulting parameter distributions (obtained using Markov Chain Monte Carlo, MCMC, methods for Bayesian inference) shown in Figure 1.
@@
However, naive application of Bayesian methods do not necessarily provide satisfactory parameter uncertainties for these models. For example, Figure 1 shows that some parameters, such as $V_{0}$, exhibit reasonable convergence toward those identified in the literature (inlay table in Figure 1), while others show wide uncertainty and do not converge toward parameters from the literature. Although this parameter estimation exercise is only preliminary, and could almost certainly be improved with informed priors from expert modelers and perhaps improved MCMC methods, it exhibits the difficulty in Bayesian model calibration – even when models have a modest number of parameters. However, many physics-based and empirical EOS models have a much larger number of parameters. For example, the Carbon EOS in the Radiative Emissivity and Opacity of Dense Plasmas (REODP) code Miloshevsky and Hassanein (2015) has 17 parameters for the diamond phase of carbon (plus 17 each for the BC8, SC, SH phases and 30 parameters for the liquid phase), which makes even deterministic calibration a massive undertaking Benedict et al. (2014). Calibrating this set of parameters using Bayesian inference is practically impossible without huge data sets and highly specialized expertise.
@@ III.2 Hugoniot derivation from the GP EOS model
The Hugoniot equation can be expressed as,
$H(V,T)=E(V,T)-E_{0}+\frac{1}{2}\left(V-V_{0}\right)\left(P(V,T)+P_{0}\right)$ (33)
$H=f\left(P,E\right)=f\left(\mu_{P},\mu_{E}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)$ (34)
Since both $P$ and $E$ are GPs, $H$ (a linear function of $P$ and $E$) is also a GP. Taking the expectation of Eq. (34) yields:
$\mathop{{}\mathbb{E}}\left[H\right]=\mathop{{}\mathbb{E}}\left[f\left(P,E\right)\right]=f\left(\mu_{P},\mu_{E}\right)=\mu_{H}$ (35)
The covariance of the $H$ GP is obtained by:
$\mathop{{}\mathbb{E}}\Big{[}(H-\mathop{{}\mathbb{E}}[H])(H-\mathop{{}\mathbb{E}}[H])\Big{]}$ $=\mathop{{}\mathbb{E}}\bigg{[}\bigg{(}\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)\bigg{)}$ (36)
The cross-covariance of the $H$ and $P$ GPs is given by
$\mathop{{}\mathbb{E}}\Big{[}(H-\mathop{{}\mathbb{E}}[H])(P-\mathop{{}\mathbb{E}}[P])\Big{]}$ $=\mathop{{}\mathbb{E}}\bigg{[}\bigg{(}\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)\bigg{)}\bigg{(}\left(P-\mu_{P}\right)\bigg{)}\bigg{]}$ (37)
$K_{HE}\left(\mathbf{X},\mathbf{X}^{\prime}\right)=K_{EE}\left(\mathbf{X},\mathbf{X}^{\prime}\right)+K_{EP}\left(\mathbf{X},\mathbf{X}^{\prime}\right)$ (38)
From Eqs. (35) and (36), the $H$ GP prior is given as
@@ III.3 Unified GP EOS model learned from multiple data sources
We can train the joint GP using any combination of available data and impose the thermodynamic constraints similar to the steps described in Section III.1 to obtain the predictive distribution of the joint GP of $P$, $E$, and $H$ at any test point $\mathbf{X}_{*}$. Further, we can condition on $H$ by first partitioning the block covariance of the joint GP in Eq. (43) as follows
$\left[\begin{array}[]{c}P\\ E\\ H\end{array}\right]\mid\mathbf{X}_{P},\mathbf{X}_{E},\mathbf{X}_{H}\sim GP\left(\left[\begin{array}[]{c}\mu_{P}\left(\mathbf{X}_{P}\right)\\ \mu_{E}\left(\mathbf{X}_{E}\right)\\ \mu_{H}\left(\mathbf{X}_{H}\right)\end{array}\right],\left[\begin{array}[]{cc|c}K_{PP}\left(\mathbf{X}_{P},\mathbf{X}_{P}\right)&K_{PE}\left(\mathbf{X}_{P},\mathbf{X}_{E}\right)&K_{PH}\left(\mathbf{X}_{P},\mathbf{X}_{H}\right)\\ K_{EP}\left(\mathbf{X}_{E},\mathbf{X}_{P}\right)&K_{EE}\left(\mathbf{X}_{E},\mathbf{X}_{E}\right)&K_{EH}\left(\mathbf{X}_{E},\mathbf{X}_{H}\right)\\ \hline\cr K_{HP}\left(\mathbf{X}_{H},\mathbf{X}_{P}\right)&K_{HE}\left(\mathbf{X}_{H},\mathbf{X}_{E}\right)&K_{HH}\left(\mathbf{X}_{H},\mathbf{X}_{H}\right)\\ \end{array}\right]\right).$ (44)
We denote this block covariance by
@@ IV.1 Constrained GP EOS model for Diamond
Figure 3: Plots of the thermodynamic stability conditions at each point on a $50\times 50$ grid of $(V,T)$. Regions of constraint violation are shaded. (a) Probabilistic stability constraint for constrained Pressure GP EOS model showing that the constraints are not violated. (b) Probabilistic stability constraint for constrained Energy GP EOS model showing that the constraints are not violated. (c) Probabilistic stability constraint for unconstrained Pressure GP EOS model showing that this specific constraint is not violated even when unconstrained. (d) Probabilistic stability constraint for unconstrained Energy GP EOS model showing that the constraint is violated at many points.
These results are supported by plotting the thermodynamic stability conditions for 2500 points from a $50\times 50$ grid of $(V,T)$ points in Figures 3(a) and (b) for the constrained P-GP and E-GP, respectively. We see that the thermodynamic stability constraints are satisfied across the domain in the proposed constrained EOS model. Meanwhile, Figures 3(c) and (d) show the stability constraints for the unconstrained P-GP and E-GP EOS models, respectively, where we see that the energy stability constraint (Eq. (32)) is violated at many point (819 points in total) in the domain. This results in a negative specific heat at these points, which will likely cause a hydrodynamics simulation to crash. The pressure stability constraint, on the other hand, is not violated even in the unconstrained GP. However, we note that the unconstrained GP EOS does not satisfy the thermodynamic consistency constraint because the covariance model is not physics informed.

### v2

@@ Abstract
Numerical simulations of high energy-density experiments require equation of state (EOS) models that relate a material’s thermodynamic state variables – specifically pressure, volume/density, energy, and temperature. EOS models are typically constructed using a semi-empirical parametric methodology, which assumes a physics-informed functional form with many tunable parameters calibrated using experimental/simulation data. Since there are inherent uncertainties in the calibration data (parametric uncertainty) and the assumed functional EOS form (model uncertainty), it is essential to perform uncertainty quantification (UQ) to improve confidence in the EOS predictions. Model uncertainty is challenging for UQ studies since it requires exploring the space of all possible physically consistent functional forms. Thus, it is often neglected in favor of parametric uncertainty, which is easier to quantify without violating thermodynamic laws. This work presents a data-driven machine learning approach to constructing EOS models that naturally captures model uncertainty while satisfying the necessary thermodynamic consistency and stability constraints. We propose a novel framework based on physics-informed Gaussian process regression (GPR) that automatically captures total uncertainty in the EOS and can be jointly trained on both simulation and experimental data sources. A GPR model for the shock Hugoniot is derived and its uncertainties are quantified using the proposed framework. We apply the proposed model to learn the EOS for the diamond solid state of carbon, using both density functional theory data and experimental shock Hugoniot data to train the model and show that the prediction uncertainty is reduced by considering the thermodynamic constraints.
@@ I Introduction
EOS models are typically constructed using semi-empirical functions where the functional form is motivated by the physics, and the parameters are calibrated using a complex combination of experimental and first-principles simulation data from a variety of sources. Once calibrated, the EOS model can be used to interpolate and extrapolate over the wide range of input states needed in hydrodynamic simulations. The semi-empirical approach is subject to two sources of uncertainty: uncertainty in the values of the parameters in the EOS model (parameter uncertainty) and uncertainty in the form of the EOS model itself (model uncertainty). Both sources must be quantified simultaneously to give a complete picture of the total EOS uncertainty. While parametric uncertainty in the EOS has been addressed in several recent works, model uncertainty remains a significant challenge. In this work, we describe a new machine learning based approach to UQ which accounts for all sources of uncertainty and provides an analytical framework for combining heterogeneous data sources into a single, uncertainty-aware EOS model.
@@ II.1 Relevant Prior Work
In the standard setting, EOS parameter calibration depends on individual modelers who leverage domain knowledge and expertise to align EOS predictions with given experimental and simulation data. Recently, it has been common to pose the calibration process as an optimization problem to solve for the best parameters that give the least prediction error compared to available dataForte et al. (2018); Bergh, Wedberg, and Lundgren (2018); Cox and Christie (2015); Myint et al. (2021). This approach can naturally be extended to capture parametric uncertainty by considering uncertainties in the calibration data sets. Ali et al. (2020) proposed a method that considers small perturbations in experimental data to calibrate model parameters using an optimization routine and propagate the experimental uncertainties using Monte Carlo simulation through the EOS models. Brown and Hund (2018) apply Bayesian model calibration to estimate parameters using dynamic material properties experiments under extreme conditions. Lindquist and Jadrich (2022) proposed a Bayesian framework to perform UQ of a multi-phase EOS model for carbon by accounting for calibration data uncertainty and yielding an ensemble of model parameters set. Walters et al. (2018) used a Bayesian statistical approach to quantify parametric uncertainty by coupling hydrocode simulations and velocimetry measurements of a series of plate impact experiments of an aluminum alloy. Robinson et al. (2013) have quantified uncertainties in the EOS stemming from the measurement noise in the experimental data and propagated these uncertainties for hydrocode simulations. They have emphasized the crucial need to integrate model uncertainty into consideration, identifying it as a significant factor.
@@ II.2 Uncertainty in Parametric EOS
UQ for existing parametric EOS models is further limited by the prescribed form of these models. Although often derived from physical principles, these models are nonetheless built upon assumptions, simplifications, and approximations of known physics, while neglecting physics that are poorly understood. Consequently, these models may be very accurate in certain regimes (e.g. of $T$ and $P$) and inadequate in others. The resulting uncertainty in these predictions is referred to as model-form uncertainty and cannot be accounted for in existing parametric models. In certain cases, competing parametric models can be compared and selected using Bayesian model selection. However, this requires the computation of the evidence term (denominator in Eq. (3)), which poses significant practical challenges. Model-form uncertainty, combined with parametric uncertainty, results in a range of outputs for a fixed input state, thus yielding an ensemble of valid EOS models. Hence, it is also necessary to quantify these model-form uncertainties in a rigorous UQ framework to enhance our confidence in EOS predictions. Several approaches exist in the literature to address the model form uncertainties in Bayesian model calibration. These include the Kennedy and O’Hagan (KOH) approachKennedy and O’Hagan (2001), the hierarchical Bayesian approachBehmanesh et al. (2015), and the Bayesian model averaging approachPark and Grandhi (2011); Hoeting et al. (1999), among others. Specifically, in the KOH approach, a discrepancy model is introduced into the mathematical model to address model form uncertainty. However, these methodologies have not been extensively explored in the context of EOS models, partly due to challenges associated with their implementation and computational complexity. Our approach is complementary to these previous works as we can use it to introduce physics constraints into the existing uncertainty frameworks as appropriate for a specific problem; for example, we could add a physically constrained GP correction term to the physics-based parametric models in the manner of KOH. Significant research efforts have been made in quantifying parametric uncertainty; however, model-form UQ for equations of state has not received adequate attention, and only a few publications are available in the literature Kamga et al. (2014); Gaffney, Yang, and Ali (2022). The biggest hurdle in quantifying model-form uncertainty is enumerating all the potential mappings that could form physics-consistent EOSs. In this work, we develop a framework using Gaussian process regression that automatically explores the range of physics-consistent EOS models to capture both sources of uncertainty. The proposed method is non-parametric and data-driven, yet satisfies both thermodynamic consistency and stability constraints.
@@ II.3 Parametric EOS with Uncertainty: An Illustration
$\theta_{D}=\theta_{D_{0}}x^{-1.5}\exp[\gamma_{1}(1-x^{3q})/q]$ (7)
@@
Fitting the model therefore requires a complicated calibration process to infer the following vector of six parameters $\alpha=\{V_{0},K_{0},K^{\prime}_{0},\theta_{0},\gamma_{1},q\}$ from data at various pressures and temperatures Dewaele et al. (2008); Occelli, Loubeyre, and LeToullec (2003). This calibration has been performed in the literature Dewaele et al. (2008) and we have further conducted a Bayesian parameter estimation using the same data with the resulting parameter distributions shown in Figure 1. These results were generated using Markov Chain Monte Carlo, MCMC, with an Affine-invariant sampler with Stretch moves Goodman and Weare (2010), implemented in our UQpy software Olivier et al. (2020); Tsapetis et al. (2023).
@@
Our Bayesian parameter estimation, with naïve priors on the parameter values, produce well-converged posterior distributions and a high-quality fit to the experimental data. Yet the Bayesian calibration is allowed to accept values in regions where the parameters have no physical meaning; for example, the posterior distributions of the $\theta_{D_{0}}$ and $q$ parameters extend to negative values (Figure 1). These parameters describe the thermal contribution to the EOS, while data are limited to a small range of parameters in the $298^{o}K\leq T\leq 900^{o}K$ region, and so it can be expected that this model will quickly develop inconsistencies when extended over a large range of parameters. While this incorrect behavior would be relatively easy to correct through a prior that limits parameters to positive values, such an approach would be very difficult for more complex models where pathologies are difficult, if not impossible, to identify a priori. This is in addition to the well-known issues with scaling MCMC to large numbers of parameters and posterior distributions with complex structure as can be expected from poorly constrained problems.
The example in Figure 1 serves to demonstrate a key difficulty with Bayesian model calibration, even when models have a modest number of parameters. Many physics-based and empirical EOS models have a much larger number of parameters which makes the parametric approach difficult and expensive. For example, the Carbon EOS in the Radiative Emissivity and Opacity of Dense Plasmas (REODP) code Miloshevsky and Hassanein (2015) has 17 parameters for the diamond phase of carbon (plus 17 each for the BC8, SC, SH phases and 30 parameters for the liquid phase), which makes even deterministic calibration a massive undertaking Benedict et al. (2014). Calibrating this set of parameters using Bayesian inference is practically impossible without huge data sets and highly specialized expertise.
@@ III.2 Hugoniot derivation from the GP EOS model
From Eqs. (A.3) and (A.4), the $H$ GP prior is given as
@@ III.3 Unified GP EOS model learned from multiple data sources
We can train the joint GP using any combination of available data and impose the thermodynamic constraints similar to the steps described in Section III.1 to obtain the predictive distribution of the joint GP of $P$, $E$, and $H$ at any test point $\mathbf{X}_{*}$. Further, we can condition on $H$ by first partitioning the block covariance of the joint GP in Eq. (37) in a similar manner as done previously. We denote this block covariance by
@@ IV.1 Constrained GP EOS model for Diamond
Figure 3: Plots of the thermodynamic stability conditions in the $(V,T)$ space. Regions of constraint violation are represented by red. (a) Probabilistic stability constraint for constrained Pressure GP EOS model showing that the constraints are not violated. (b) Probabilistic stability constraint for constrained Energy GP EOS model showing that the constraints are not violated. (c) Probabilistic stability constraint for unconstrained Pressure GP EOS model showing that this specific constraint is not violated even when unconstrained. (d) Probabilistic stability constraint for unconstrained Energy GP EOS model showing that the constraint is violated in regions characterized by high temperatures across volumes within the domain.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 4: Plots of the validation study between the proposed GP EOS model and parametric Benedict EOS model:Benedict et al. (2014) (a) Percentage $*{}^{*}$ between mean pressure GP EOS and Benedict pressure EOS. (b) Percentage $*{}^{*}$ between mean energy GP EOS and Benedict energy EOS. (c) Comparison of isothermal bulk modulus ($K_{T}$) obtained from GP EOS model (with $95\%$ confidence bounds) and Benedict EOSBenedict et al. (2014) for $T_{1}=2490K$ and $T_{2}=7510K$. (d) Comparison of specific isochoric heat capacity ($c_{V}$) obtained from GP EOS model (with $95\%$ confidence bounds) and Benedict EOSBenedict et al. (2014) for $V_{1}=4.8\ A^{3}/atom$, $V_{2}=4.2\ A^{3}/atom$ and $V_{3}=3.6\ A^{3}/atom$. *$\Big{|}\frac{X_{\text{mean GP EOS}}-X_{\text{Benedict EOS}}}{X_{\text{BenedictEOS}}}\Big{|}\times 100$
These results are supported by plotting the thermodynamic stability conditions in the $(V,T)$ space in Figures 3(a) and (b) for the constrained P-GP and E-GP, respectively. We see that the thermodynamic stability constraints are satisfied across the domain in the proposed constrained EOS model. Meanwhile, Figures 3(c) and (d) show the stability constraints for the unconstrained P-GP and E-GP EOS models, respectively, where we see that the energy stability constraint (Eq. (32)) is violated in regions characterized by high temperatures across volumes within the domain.. This results in a negative specific heat in these regions, which will likely cause a hydrodynamics simulation to crash. The pressure stability constraint, on the other hand, is not violated even in the unconstrained GP. However, we note that the unconstrained GP EOS does not satisfy the thermodynamic consistency constraint because the covariance model is not physics informed.
@@
We have validated the results obtained from our data driven GP EOS model with the parameteric Benedict EOSBenedict et al. (2014) in Figure 4. In Figures 4(a) and (b), we first plotted the percentage error for pressure and energy between the ones predicted by the Benedict EOS Benedict et al. (2014) and the mean of our GP EOS. As can be observed from the Figure 4(a), the percentage error in pressure remains relatively low across most regions in the $(V,\ T)$ space. In specific regions characterized by high volume and low temperature, as well as high volume and temperature, there is a moderate deviation, typically ranging between $5-10\%$. For these regions, our pressure GP EOS also gives the highest uncertainty. Similarly, in Figure 4(b), the percentage error in energy is relatively low across most $(V,\ T)$ regions with a moderate deviation (around $3-5.5\%$) for high volume and temperature. Again, our energy GP EOS gives the highest uncertainty for this region. This indicates the effectiveness of our model in accurately representing the characteristic uncertainty in the model.
We also obtained the specific isochoric heat capacity ($c_{V}$) and isothermal bulk modulus ($K_{T}$) from the Benedict EOS Benedict et al. (2014) and the proposed GP EOS model. These quantities, which were not explicitly included in the training of our model but are derived from it by numerical differentiation, serve to further validate the GP EOS and the uncertainties from it. In Figure 4(c), we plot the isothermal Bulk modulus against volume. The figure shows that the mean $K_{T}$ obtained from our model aligns well with the values derived from the Benedict EOS Benedict et al. (2014). Any observed deviation is effectively characterized within the $95\%$ confidence bounds. Additionally, it can be seen that $K_{T}$ does not vary appreciably with temperature within the specified temperature range.
Similarly, for $c_{V}$ we can observe good agreement between the mean $c_{V}$ obtained from our model and Benedict EOS in Figure 4(d). As pointed out in the literature Benedict et al. (2014); Correa et al. (2008), surprisingly, the specific heat is observed to be indistinguishable from $3k_{B}/atom$ up to 20,000 K and is also very nearly independent of volume. Our model captures these observations within the 95% confidence intervals, but suggests there is significant uncertainty in $c_{V}$ particularly at high temperature. This uncertainty results from seemingly small uncertainties in the EOS that grow considerably when differentiated to compute $c_{V}=\dfrac{\partial E}{\partial T}\biggr{|}_{V}$.